%merlin.mbs apsrev4-1.bst 2010-07-25 4.21a (PWD, AO, DPC) hacked
%Control: key (0)
%Control: author (8) initials jnrlst
%Control: editor formatted (1) identically to author
%Control: production of article title (-1) disabled
%Control: page (0) single
%Control: year (1) truncated
%Control: production of eprint (0) enabled
\begin{thebibliography}{36}%
\makeatletter
\providecommand \@ifxundefined [1]{%
 \@ifx{#1\undefined}
}%
\providecommand \@ifnum [1]{%
 \ifnum #1\expandafter \@firstoftwo
 \else \expandafter \@secondoftwo
 \fi
}%
\providecommand \@ifx [1]{%
 \ifx #1\expandafter \@firstoftwo
 \else \expandafter \@secondoftwo
 \fi
}%
\providecommand \natexlab [1]{#1}%
\providecommand \enquote  [1]{``#1''}%
\providecommand \bibnamefont  [1]{#1}%
\providecommand \bibfnamefont [1]{#1}%
\providecommand \citenamefont [1]{#1}%
\providecommand \href@noop [0]{\@secondoftwo}%
\providecommand \href [0]{\begingroup \@sanitize@url \@href}%
\providecommand \@href[1]{\@@startlink{#1}\@@href}%
\providecommand \@@href[1]{\endgroup#1\@@endlink}%
\providecommand \@sanitize@url [0]{\catcode `\\12\catcode `\$12\catcode
  `\&12\catcode `\#12\catcode `\^12\catcode `\_12\catcode `\%12\relax}%
\providecommand \@@startlink[1]{}%
\providecommand \@@endlink[0]{}%
\providecommand \url  [0]{\begingroup\@sanitize@url \@url }%
\providecommand \@url [1]{\endgroup\@href {#1}{\urlprefix }}%
\providecommand \urlprefix  [0]{URL }%
\providecommand \Eprint [0]{\href }%
\providecommand \doibase [0]{http://dx.doi.org/}%
\providecommand \selectlanguage [0]{\@gobble}%
\providecommand \bibinfo  [0]{\@secondoftwo}%
\providecommand \bibfield  [0]{\@secondoftwo}%
\providecommand \translation [1]{[#1]}%
\providecommand \BibitemOpen [0]{}%
\providecommand \bibitemStop [0]{}%
\providecommand \bibitemNoStop [0]{.\EOS\space}%
\providecommand \EOS [0]{\spacefactor3000\relax}%
\providecommand \BibitemShut  [1]{\csname bibitem#1\endcsname}%
\let\auto@bib@innerbib\@empty
%</preamble>
\bibitem [{\citenamefont {Nir}\ and\ \citenamefont {Quinn}(1991)}]{Nir:1991cu}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Y.}~\bibnamefont
  {Nir}}\ and\ \bibinfo {author} {\bibfnamefont {H.~R.}\ \bibnamefont
  {Quinn}},\ }\href {\doibase 10.1103/PhysRevLett.67.541} {\bibfield  {journal}
  {\bibinfo  {journal} {Phys. Rev. Lett.}\ }\textbf {\bibinfo {volume} {67}},\
  \bibinfo {pages} {541} (\bibinfo {year} {1991})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Lipkin}\ \emph {et~al.}(1991)\citenamefont {Lipkin},
  \citenamefont {Nir}, \citenamefont {Quinn},\ and\ \citenamefont
  {Snyder}}]{Lipkin:1991st}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {H.~J.}\ \bibnamefont
  {Lipkin}}, \bibinfo {author} {\bibfnamefont {Y.}~\bibnamefont {Nir}},
  \bibinfo {author} {\bibfnamefont {H.~R.}\ \bibnamefont {Quinn}}, \ and\
  \bibinfo {author} {\bibfnamefont {A.}~\bibnamefont {Snyder}},\ }\href
  {\doibase 10.1103/PhysRevD.44.1454} {\bibfield  {journal} {\bibinfo
  {journal} {Phys. Rev. D}\ }\textbf {\bibinfo {volume} {44}},\ \bibinfo
  {pages} {1454} (\bibinfo {year} {1991})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Gronau}(1991)}]{Gronau:1991dq}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {M.}~\bibnamefont
  {Gronau}},\ }\href {\doibase 10.1016/0370-2693(91)90071-W} {\bibfield
  {journal} {\bibinfo  {journal} {Phys. Lett. B}\ }\textbf {\bibinfo {volume}
  {265}},\ \bibinfo {pages} {389} (\bibinfo {year} {1991})}\BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Gronau}(2005)}]{Gronau:2005kz}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {M.}~\bibnamefont
  {Gronau}},\ }\href {\doibase 10.1016/j.physletb.2005.09.014} {\bibfield
  {journal} {\bibinfo  {journal} {Phys. Lett. B}\ }\textbf {\bibinfo {volume}
  {627}},\ \bibinfo {pages} {82} (\bibinfo {year} {2005})},\ \Eprint
  {http://arxiv.org/abs/hep-ph/0508047} {arXiv:hep-ph/0508047} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Gronau}\ and\ \citenamefont
  {Rosner}(1999)}]{Gronau:1998ep}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {M.}~\bibnamefont
  {Gronau}}\ and\ \bibinfo {author} {\bibfnamefont {J.~L.}\ \bibnamefont
  {Rosner}},\ }\href {\doibase 10.1103/PhysRevD.59.113002} {\bibfield
  {journal} {\bibinfo  {journal} {Phys. Rev. D}\ }\textbf {\bibinfo {volume}
  {59}},\ \bibinfo {pages} {113002} (\bibinfo {year} {1999})},\ \Eprint
  {http://arxiv.org/abs/hep-ph/9809384} {arXiv:hep-ph/9809384} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Duh}\ \emph {et~al.}(2013)\citenamefont {Duh} \emph
  {et~al.}}]{Belle:2012dmz}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Y.~T.}\ \bibnamefont
  {Duh}} \emph {et~al.} (\bibinfo {collaboration} {Belle}),\ }\href {\doibase
  10.1103/PhysRevD.87.031103} {\bibfield  {journal} {\bibinfo  {journal} {Phys.
  Rev. D}\ }\textbf {\bibinfo {volume} {87}},\ \bibinfo {pages} {031103}
  (\bibinfo {year} {2013})},\ \Eprint {http://arxiv.org/abs/1210.1348}
  {arXiv:1210.1348 [hep-ex]} \BibitemShut {NoStop}%
\bibitem [{\citenamefont {Lees}\ \emph {et~al.}(2013)\citenamefont {Lees} \emph
  {et~al.}}]{BaBar:2012fgk}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {J.~P.}\ \bibnamefont
  {Lees}} \emph {et~al.} (\bibinfo {collaboration} {BaBar}),\ }\href {\doibase
  10.1103/PhysRevD.87.052009} {\bibfield  {journal} {\bibinfo  {journal} {Phys.
  Rev. D}\ }\textbf {\bibinfo {volume} {87}},\ \bibinfo {pages} {052009}
  (\bibinfo {year} {2013})},\ \Eprint {http://arxiv.org/abs/1206.3525}
  {arXiv:1206.3525 [hep-ex]} \BibitemShut {NoStop}%
\bibitem [{\citenamefont {Aaltonen}\ \emph {et~al.}(2014)\citenamefont
  {Aaltonen} \emph {et~al.}}]{CDF:2014pzb}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {T.~A.}\ \bibnamefont
  {Aaltonen}} \emph {et~al.} (\bibinfo {collaboration} {CDF}),\ }\href
  {\doibase 10.1103/PhysRevLett.113.242001} {\bibfield  {journal} {\bibinfo
  {journal} {Phys. Rev. Lett.}\ }\textbf {\bibinfo {volume} {113}},\ \bibinfo
  {pages} {242001} (\bibinfo {year} {2014})},\ \Eprint
  {http://arxiv.org/abs/1403.5586} {arXiv:1403.5586 [hep-ex]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Aaij}\ \emph {et~al.}(2018)\citenamefont {Aaij} \emph
  {et~al.}}]{LHCb:2018pff}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {R.}~\bibnamefont
  {Aaij}} \emph {et~al.} (\bibinfo {collaboration} {LHCb}),\ }\href {\doibase
  10.1103/PhysRevD.98.032004} {\bibfield  {journal} {\bibinfo  {journal} {Phys.
  Rev. D}\ }\textbf {\bibinfo {volume} {98}},\ \bibinfo {pages} {032004}
  (\bibinfo {year} {2018})},\ \Eprint {http://arxiv.org/abs/1805.06759}
  {arXiv:1805.06759 [hep-ex]} \BibitemShut {NoStop}%
\bibitem [{\citenamefont {Amhis}\ \emph {et~al.}(2021)\citenamefont {Amhis}
  \emph {et~al.}}]{HFLAV:2019otj}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Y.~S.}\ \bibnamefont
  {Amhis}} \emph {et~al.} (\bibinfo {collaboration} {HFLAV}),\ }\href {\doibase
  10.1140/epjc/s10052-020-8156-7} {\bibfield  {journal} {\bibinfo  {journal}
  {Eur. Phys. J. C}\ }\textbf {\bibinfo {volume} {81}},\ \bibinfo {pages} {226}
  (\bibinfo {year} {2021})},\ \Eprint {http://arxiv.org/abs/1909.12524}
  {arXiv:1909.12524 [hep-ex]} \BibitemShut {NoStop}%
\bibitem [{\citenamefont {Aaij}\ \emph {et~al.}(2021)\citenamefont {Aaij} \emph
  {et~al.}}]{LHCb:2020dpr}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {R.}~\bibnamefont
  {Aaij}} \emph {et~al.} (\bibinfo {collaboration} {LHCb}),\ }\href {\doibase
  10.1103/PhysRevLett.126.091802} {\bibfield  {journal} {\bibinfo  {journal}
  {Phys. Rev. Lett.}\ }\textbf {\bibinfo {volume} {126}},\ \bibinfo {pages}
  {091802} (\bibinfo {year} {2021})},\ \Eprint
  {http://arxiv.org/abs/2012.12789} {arXiv:2012.12789 [hep-ex]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Buras}\ \emph {et~al.}(2003)\citenamefont {Buras},
  \citenamefont {Fleischer}, \citenamefont {Recksiegel},\ and\ \citenamefont
  {Schwab}}]{Buras:2003yc}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {A.~J.}\ \bibnamefont
  {Buras}}, \bibinfo {author} {\bibfnamefont {R.}~\bibnamefont {Fleischer}},
  \bibinfo {author} {\bibfnamefont {S.}~\bibnamefont {Recksiegel}}, \ and\
  \bibinfo {author} {\bibfnamefont {F.}~\bibnamefont {Schwab}},\ }\href
  {\doibase 10.1140/epjc/s2003-01379-9} {\bibfield  {journal} {\bibinfo
  {journal} {Eur. Phys. J. C}\ }\textbf {\bibinfo {volume} {32}},\ \bibinfo
  {pages} {45} (\bibinfo {year} {2003})},\ \Eprint
  {http://arxiv.org/abs/hep-ph/0309012} {arXiv:hep-ph/0309012} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Buras}\ \emph
  {et~al.}(2004{\natexlab{a}})\citenamefont {Buras}, \citenamefont {Fleischer},
  \citenamefont {Recksiegel},\ and\ \citenamefont {Schwab}}]{Buras:2003dj}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {A.~J.}\ \bibnamefont
  {Buras}}, \bibinfo {author} {\bibfnamefont {R.}~\bibnamefont {Fleischer}},
  \bibinfo {author} {\bibfnamefont {S.}~\bibnamefont {Recksiegel}}, \ and\
  \bibinfo {author} {\bibfnamefont {F.}~\bibnamefont {Schwab}},\ }\href
  {\doibase 10.1103/PhysRevLett.92.101804} {\bibfield  {journal} {\bibinfo
  {journal} {Phys. Rev. Lett.}\ }\textbf {\bibinfo {volume} {92}},\ \bibinfo
  {pages} {101804} (\bibinfo {year} {2004}{\natexlab{a}})},\ \Eprint
  {http://arxiv.org/abs/hep-ph/0312259} {arXiv:hep-ph/0312259} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Buras}\ \emph
  {et~al.}(2004{\natexlab{b}})\citenamefont {Buras}, \citenamefont {Fleischer},
  \citenamefont {Recksiegel},\ and\ \citenamefont {Schwab}}]{Buras:2004ub}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {A.~J.}\ \bibnamefont
  {Buras}}, \bibinfo {author} {\bibfnamefont {R.}~\bibnamefont {Fleischer}},
  \bibinfo {author} {\bibfnamefont {S.}~\bibnamefont {Recksiegel}}, \ and\
  \bibinfo {author} {\bibfnamefont {F.}~\bibnamefont {Schwab}},\ }\href
  {\doibase 10.1016/j.nuclphysb.2004.07.009} {\bibfield  {journal} {\bibinfo
  {journal} {Nucl. Phys. B}\ }\textbf {\bibinfo {volume} {697}},\ \bibinfo
  {pages} {133} (\bibinfo {year} {2004}{\natexlab{b}})},\ \Eprint
  {http://arxiv.org/abs/hep-ph/0402112} {arXiv:hep-ph/0402112} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Barger}\ \emph {et~al.}(2004)\citenamefont {Barger},
  \citenamefont {Chiang}, \citenamefont {Langacker},\ and\ \citenamefont
  {Lee}}]{Barger:2004hn}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {V.}~\bibnamefont
  {Barger}}, \bibinfo {author} {\bibfnamefont {C.-W.}\ \bibnamefont {Chiang}},
  \bibinfo {author} {\bibfnamefont {P.}~\bibnamefont {Langacker}}, \ and\
  \bibinfo {author} {\bibfnamefont {H.-S.}\ \bibnamefont {Lee}},\ }\href
  {\doibase 10.1016/j.physletb.2004.07.057} {\bibfield  {journal} {\bibinfo
  {journal} {Phys. Lett. B}\ }\textbf {\bibinfo {volume} {598}},\ \bibinfo
  {pages} {218} (\bibinfo {year} {2004})},\ \Eprint
  {http://arxiv.org/abs/hep-ph/0406126} {arXiv:hep-ph/0406126} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Baek}\ \emph {et~al.}(2005)\citenamefont {Baek},
  \citenamefont {Hamel}, \citenamefont {London}, \citenamefont {Datta},\ and\
  \citenamefont {Suprun}}]{Baek:2004rp}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {S.}~\bibnamefont
  {Baek}}, \bibinfo {author} {\bibfnamefont {P.}~\bibnamefont {Hamel}},
  \bibinfo {author} {\bibfnamefont {D.}~\bibnamefont {London}}, \bibinfo
  {author} {\bibfnamefont {A.}~\bibnamefont {Datta}}, \ and\ \bibinfo {author}
  {\bibfnamefont {D.~A.}\ \bibnamefont {Suprun}},\ }\href {\doibase
  10.1103/PhysRevD.71.057502} {\bibfield  {journal} {\bibinfo  {journal} {Phys.
  Rev. D}\ }\textbf {\bibinfo {volume} {71}},\ \bibinfo {pages} {057502}
  (\bibinfo {year} {2005})},\ \Eprint {http://arxiv.org/abs/hep-ph/0412086}
  {arXiv:hep-ph/0412086} \BibitemShut {NoStop}%
\bibitem [{\citenamefont {Chang}\ \emph {et~al.}(2009)\citenamefont {Chang},
  \citenamefont {Li},\ and\ \citenamefont {Yang}}]{Chang:2009wt}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Q.}~\bibnamefont
  {Chang}}, \bibinfo {author} {\bibfnamefont {X.-Q.}\ \bibnamefont {Li}}, \
  and\ \bibinfo {author} {\bibfnamefont {Y.-D.}\ \bibnamefont {Yang}},\ }\href
  {\doibase 10.1088/1126-6708/2009/05/056} {\bibfield  {journal} {\bibinfo
  {journal} {JHEP}\ }\textbf {\bibinfo {volume} {05}},\ \bibinfo {pages} {056}
  (\bibinfo {year} {2009})},\ \Eprint {http://arxiv.org/abs/0903.0275}
  {arXiv:0903.0275 [hep-ph]} \BibitemShut {NoStop}%
\bibitem [{\citenamefont {Fleischer}\ \emph {et~al.}(2018)\citenamefont
  {Fleischer}, \citenamefont {Jaarsma},\ and\ \citenamefont
  {Vos}}]{Fleischer:2017vrb}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {R.}~\bibnamefont
  {Fleischer}}, \bibinfo {author} {\bibfnamefont {R.}~\bibnamefont {Jaarsma}},
  \ and\ \bibinfo {author} {\bibfnamefont {K.~K.}\ \bibnamefont {Vos}},\ }\href
  {\doibase 10.1016/j.physletb.2018.09.015} {\bibfield  {journal} {\bibinfo
  {journal} {Phys. Lett. B}\ }\textbf {\bibinfo {volume} {785}},\ \bibinfo
  {pages} {525} (\bibinfo {year} {2018})},\ \Eprint
  {http://arxiv.org/abs/1712.02323} {arXiv:1712.02323 [hep-ph]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Beaudry}\ \emph {et~al.}(2018)\citenamefont
  {Beaudry}, \citenamefont {Datta}, \citenamefont {London}, \citenamefont
  {Rashed},\ and\ \citenamefont {Roux}}]{Beaudry:2017gtw}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {N.~B.}\ \bibnamefont
  {Beaudry}}, \bibinfo {author} {\bibfnamefont {A.}~\bibnamefont {Datta}},
  \bibinfo {author} {\bibfnamefont {D.}~\bibnamefont {London}}, \bibinfo
  {author} {\bibfnamefont {A.}~\bibnamefont {Rashed}}, \ and\ \bibinfo {author}
  {\bibfnamefont {J.-S.}\ \bibnamefont {Roux}},\ }\href {\doibase
  10.1007/JHEP01(2018)074} {\bibfield  {journal} {\bibinfo  {journal} {JHEP}\
  }\textbf {\bibinfo {volume} {01}},\ \bibinfo {pages} {074} (\bibinfo {year}
  {2018})},\ \Eprint {http://arxiv.org/abs/1709.07142} {arXiv:1709.07142
  [hep-ph]} \BibitemShut {NoStop}%
\bibitem [{\citenamefont {Bhattacharya}\ \emph {et~al.}(2021)\citenamefont
  {Bhattacharya}, \citenamefont {Datta}, \citenamefont {Marfatia},
  \citenamefont {Nandi},\ and\ \citenamefont {Waite}}]{Bhattacharya:2021shk}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {B.}~\bibnamefont
  {Bhattacharya}}, \bibinfo {author} {\bibfnamefont {A.}~\bibnamefont {Datta}},
  \bibinfo {author} {\bibfnamefont {D.}~\bibnamefont {Marfatia}}, \bibinfo
  {author} {\bibfnamefont {S.}~\bibnamefont {Nandi}}, \ and\ \bibinfo {author}
  {\bibfnamefont {J.}~\bibnamefont {Waite}},\ }\href {\doibase
  10.1103/PhysRevD.104.L051701} {\bibfield  {journal} {\bibinfo  {journal}
  {Phys. Rev. D}\ }\textbf {\bibinfo {volume} {104}},\ \bibinfo {pages}
  {051701} (\bibinfo {year} {2021})},\ \Eprint
  {http://arxiv.org/abs/2104.03947} {arXiv:2104.03947 [hep-ph]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Bauer}\ \emph {et~al.}(2006)\citenamefont {Bauer},
  \citenamefont {Rothstein},\ and\ \citenamefont {Stewart}}]{Bauer:2005kd}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {C.~W.}\ \bibnamefont
  {Bauer}}, \bibinfo {author} {\bibfnamefont {I.~Z.}\ \bibnamefont
  {Rothstein}}, \ and\ \bibinfo {author} {\bibfnamefont {I.~W.}\ \bibnamefont
  {Stewart}},\ }\href {\doibase 10.1103/PhysRevD.74.034010} {\bibfield
  {journal} {\bibinfo  {journal} {Phys. Rev. D}\ }\textbf {\bibinfo {volume}
  {74}},\ \bibinfo {pages} {034010} (\bibinfo {year} {2006})},\ \Eprint
  {http://arxiv.org/abs/hep-ph/0510241} {arXiv:hep-ph/0510241} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Cheng}\ and\ \citenamefont
  {Chua}(2009)}]{Cheng:2009cn}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {H.-Y.}\ \bibnamefont
  {Cheng}}\ and\ \bibinfo {author} {\bibfnamefont {C.-K.}\ \bibnamefont
  {Chua}},\ }\href {\doibase 10.1103/PhysRevD.80.114008} {\bibfield  {journal}
  {\bibinfo  {journal} {Phys. Rev. D}\ }\textbf {\bibinfo {volume} {80}},\
  \bibinfo {pages} {114008} (\bibinfo {year} {2009})},\ \Eprint
  {http://arxiv.org/abs/0909.5229} {arXiv:0909.5229 [hep-ph]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Cheng}\ \emph {et~al.}(2015)\citenamefont {Cheng},
  \citenamefont {Chiang},\ and\ \citenamefont {Kuo}}]{Cheng:2014rfa}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {H.-Y.}\ \bibnamefont
  {Cheng}}, \bibinfo {author} {\bibfnamefont {C.-W.}\ \bibnamefont {Chiang}}, \
  and\ \bibinfo {author} {\bibfnamefont {A.-L.}\ \bibnamefont {Kuo}},\ }\href
  {\doibase 10.1103/PhysRevD.91.014011} {\bibfield  {journal} {\bibinfo
  {journal} {Phys. Rev. D}\ }\textbf {\bibinfo {volume} {91}},\ \bibinfo
  {pages} {014011} (\bibinfo {year} {2015})},\ \Eprint
  {http://arxiv.org/abs/1409.5026} {arXiv:1409.5026 [hep-ph]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Zhou}\ \emph {et~al.}(2017)\citenamefont {Zhou},
  \citenamefont {Zhang}, \citenamefont {Lyu},\ and\ \citenamefont
  {L\"u}}]{Zhou:2016jkv}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {S.-H.}\ \bibnamefont
  {Zhou}}, \bibinfo {author} {\bibfnamefont {Q.-A.}\ \bibnamefont {Zhang}},
  \bibinfo {author} {\bibfnamefont {W.-R.}\ \bibnamefont {Lyu}}, \ and\
  \bibinfo {author} {\bibfnamefont {C.-D.}\ \bibnamefont {L\"u}},\ }\href
  {\doibase 10.1140/epjc/s10052-017-4685-0} {\bibfield  {journal} {\bibinfo
  {journal} {Eur. Phys. J. C}\ }\textbf {\bibinfo {volume} {77}},\ \bibinfo
  {pages} {125} (\bibinfo {year} {2017})},\ \Eprint
  {http://arxiv.org/abs/1608.02819} {arXiv:1608.02819 [hep-ph]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Kundu}\ \emph {et~al.}(2021)\citenamefont {Kundu},
  \citenamefont {Patra},\ and\ \citenamefont {Roy}}]{Kundu:2021emt}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {A.}~\bibnamefont
  {Kundu}}, \bibinfo {author} {\bibfnamefont {S.~K.}\ \bibnamefont {Patra}}, \
  and\ \bibinfo {author} {\bibfnamefont {S.}~\bibnamefont {Roy}},\ }\href
  {\doibase 10.1103/PhysRevD.104.095025} {\bibfield  {journal} {\bibinfo
  {journal} {Phys. Rev. D}\ }\textbf {\bibinfo {volume} {104}},\ \bibinfo
  {pages} {095025} (\bibinfo {year} {2021})},\ \Eprint
  {http://arxiv.org/abs/2106.15633} {arXiv:2106.15633 [hep-ph]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Gross}\ \emph {et~al.}(1979)\citenamefont {Gross},
  \citenamefont {Treiman},\ and\ \citenamefont {Wilczek}}]{Gross:1979ur}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {D.~J.}\ \bibnamefont
  {Gross}}, \bibinfo {author} {\bibfnamefont {S.~B.}\ \bibnamefont {Treiman}},
  \ and\ \bibinfo {author} {\bibfnamefont {F.}~\bibnamefont {Wilczek}},\ }\href
  {\doibase 10.1103/PhysRevD.19.2188} {\bibfield  {journal} {\bibinfo
  {journal} {Phys. Rev. D}\ }\textbf {\bibinfo {volume} {19}},\ \bibinfo
  {pages} {2188} (\bibinfo {year} {1979})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Leutwyler}(1996)}]{Leutwyler:1996np}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {H.}~\bibnamefont
  {Leutwyler}},\ }\href {\doibase 10.1016/0370-2693(96)00167-0} {\bibfield
  {journal} {\bibinfo  {journal} {Phys. Lett. B}\ }\textbf {\bibinfo {volume}
  {374}},\ \bibinfo {pages} {181} (\bibinfo {year} {1996})},\ \Eprint
  {http://arxiv.org/abs/hep-ph/9601236} {arXiv:hep-ph/9601236} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Feldmann}\ \emph {et~al.}(1999)\citenamefont
  {Feldmann}, \citenamefont {Kroll},\ and\ \citenamefont
  {Stech}}]{Feldmann:1998sh}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {T.}~\bibnamefont
  {Feldmann}}, \bibinfo {author} {\bibfnamefont {P.}~\bibnamefont {Kroll}}, \
  and\ \bibinfo {author} {\bibfnamefont {B.}~\bibnamefont {Stech}},\ }\href
  {\doibase 10.1016/S0370-2693(99)00085-4} {\bibfield  {journal} {\bibinfo
  {journal} {Phys. Lett. B}\ }\textbf {\bibinfo {volume} {449}},\ \bibinfo
  {pages} {339} (\bibinfo {year} {1999})},\ \Eprint
  {http://arxiv.org/abs/hep-ph/9812269} {arXiv:hep-ph/9812269} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Escribano}\ and\ \citenamefont
  {Royo}(2020)}]{Escribano:2020jdy}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {R.}~\bibnamefont
  {Escribano}}\ and\ \bibinfo {author} {\bibfnamefont {E.}~\bibnamefont
  {Royo}},\ }\href {\doibase 10.1016/j.physletb.2020.135534} {\bibfield
  {journal} {\bibinfo  {journal} {Phys. Lett. B}\ }\textbf {\bibinfo {volume}
  {807}},\ \bibinfo {pages} {135534} (\bibinfo {year} {2020})},\ \Eprint
  {http://arxiv.org/abs/2003.08379} {arXiv:2003.08379 [hep-ph]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Kordov}\ \emph {et~al.}(2021)\citenamefont {Kordov}
  \emph {et~al.}}]{Kordov:2021eqx}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Z.~R.}\ \bibnamefont
  {Kordov}} \emph {et~al.},\ }\href@noop {} {\  (\bibinfo {year} {2021})},\
  \Eprint {http://arxiv.org/abs/2110.11533} {arXiv:2110.11533 [hep-lat]}
  \BibitemShut {NoStop}%
\bibitem [{\citenamefont {Gronau}\ \emph {et~al.}(2006)\citenamefont {Gronau},
  \citenamefont {Grossman}, \citenamefont {Raz},\ and\ \citenamefont
  {Rosner}}]{Gronau:2006eb}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {M.}~\bibnamefont
  {Gronau}}, \bibinfo {author} {\bibfnamefont {Y.}~\bibnamefont {Grossman}},
  \bibinfo {author} {\bibfnamefont {G.}~\bibnamefont {Raz}}, \ and\ \bibinfo
  {author} {\bibfnamefont {J.~L.}\ \bibnamefont {Rosner}},\ }\href {\doibase
  10.1016/j.physletb.2006.02.066} {\bibfield  {journal} {\bibinfo  {journal}
  {Phys. Lett. B}\ }\textbf {\bibinfo {volume} {635}},\ \bibinfo {pages} {207}
  (\bibinfo {year} {2006})},\ \Eprint {http://arxiv.org/abs/hep-ph/0601129}
  {arXiv:hep-ph/0601129} \BibitemShut {NoStop}%
\bibitem [{\citenamefont {Zhang}\ \emph {et~al.}(2020)\citenamefont {Zhang},
  \citenamefont {Yang},\ and\ \citenamefont {Guo}}]{Zhang:2020rol}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Z.-H.}\ \bibnamefont
  {Zhang}}, \bibinfo {author} {\bibfnamefont {S.}~\bibnamefont {Yang}}, \ and\
  \bibinfo {author} {\bibfnamefont {X.-H.}\ \bibnamefont {Guo}},\ }\href
  {\doibase 10.1007/JHEP10(2020)020} {\bibfield  {journal} {\bibinfo  {journal}
  {JHEP}\ }\textbf {\bibinfo {volume} {10}},\ \bibinfo {pages} {020} (\bibinfo
  {year} {2020})},\ \Eprint {http://arxiv.org/abs/2005.09157} {arXiv:2005.09157
  [hep-ph]} \BibitemShut {NoStop}%
\bibitem [{Note1()}]{Note1}%
  \BibitemOpen
  \bibinfo {note} {An extra factor $\left (1-{A_{CP}^{K^+\pi _3 }}^2\right )$
  in front of $\Delta _{\protect \REV@text {IB}}^{K^+\pi ^0 }$ is omitted as it
  is numerically very close to one.}\BibitemShut {Stop}%
\bibitem [{Note2()}]{Note2}%
  \BibitemOpen
  \bibinfo {note} {Contributions from the branching ratios, which are
  proportional to $ A_{CP}^{K^{+,0}\pi ^0} \epsilon ^{(\prime )}$, have been
  neglected in $\Delta _{\protect \REV@text {IB}}$.}\BibitemShut {Stop}%
\bibitem [{\citenamefont {Kroll}(2005)}]{Kroll:2005sd}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {P.}~\bibnamefont
  {Kroll}},\ }\href {\doibase 10.1142/S0217732305018633} {\bibfield  {journal}
  {\bibinfo  {journal} {Mod. Phys. Lett. A}\ }\textbf {\bibinfo {volume}
  {20}},\ \bibinfo {pages} {2667} (\bibinfo {year} {2005})},\ \Eprint
  {http://arxiv.org/abs/hep-ph/0509031} {arXiv:hep-ph/0509031} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Benayoun}\ \emph {et~al.}(2021)\citenamefont
  {Benayoun}, \citenamefont {DelBuono},\ and\ \citenamefont
  {Jegerlehner}}]{Benayoun:2021ody}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {M.}~\bibnamefont
  {Benayoun}}, \bibinfo {author} {\bibfnamefont {L.}~\bibnamefont {DelBuono}},
  \ and\ \bibinfo {author} {\bibfnamefont {F.}~\bibnamefont {Jegerlehner}},\
  }\href@noop {} {\  (\bibinfo {year} {2021})},\ \Eprint
  {http://arxiv.org/abs/2105.13018} {arXiv:2105.13018 [hep-ph]} \BibitemShut
  {NoStop}%
\end{thebibliography}%

\documentclass[aps,preprint,groupedaddress,showpacs]{revtex4-1}
\usepackage{amssymb}
\usepackage[breaklinks,colorlinks,citecolor=green,urlcolor=blue,linkcolor=red]{hyperref}

\begin{document}


\title{The pollution to the $K\pi$-puzzle from the isospin-breaking $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effect}


\author{Zhen-Hua Zhang}
\affiliation{College of Nuclear Science and Technology, University of South China, Hengyang, 421001, Hunan, China}


\date{\today}

\begin{abstract}
The influence of the isospin-breaking $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effect on the $CP$-asymmetries of $B\to K\pi$ processes is examined for the first time.
It is found that this mixing effect brings a large uncertainty both to the $CP$-asymmetry sum rule of $B\to K\pi$ processes and the $CP$-asymmetry difference of the $B^+\to K^+\pi^0$ and $B^0\to K^+\pi^-$, obscuring the significance of the $K\pi$-puzzle.
This uncertainty can be so large that it is even possible to explain the $K\pi$-puzzle by the $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effect {\it alone}.
\end{abstract}


\maketitle


\section{Brief introduction to the $K \pi$-puzzle}

It is believed that the decay processes $B\to K\pi$ are good probes for new physics (NP) beyond the Standard Model (SM), as the tree-level amplitudes are suppressed, making them more sensitive to the potentially NP contributions.
Based on the isospin consideration, the amplitudes of the aforementioned weak decays $B\to K\pi$ are related through \cite{Nir:1991cu,Lipkin:1991st,Gronau:1991dq}
\begin{equation}
  \mathcal{A}_{B^+\to K^0\pi^+}+\sqrt{2}\mathcal{A}_{B^+\to K^+\pi^0}=\mathcal{A}_{{B^0}\to K^+\pi^-}+\sqrt{2}\mathcal{A}_{{B^0}\to K^0\pi^0},
\end{equation}
from which a sum rule between $CP$ asymmetries of $B\to K\pi$ processes can be deduced \cite{Gronau:2005kz}:
\begin{equation}\label{SumRule}
A_{CP}^{K^+\pi^-}+\frac{\mathcal{B}(K^0\pi^+)}{\mathcal{B}(K^+\pi^-)}\frac{\tau_0}{\tau_+}A_{CP}^{K^0\pi^+} =\frac{2\mathcal{B}(K^+\pi^0)}{\mathcal{B}(K^+\pi^-)}\frac{\tau_0}{\tau_+}A_{CP}^{K^+\pi^0}+\frac{2\mathcal{B}(K^0\pi^0)}{\mathcal{B}(K^+\pi^-)}\frac{\tau_0}{\tau_+}A_{CP}^{K^0\pi^0}.
\end{equation}
This $CP$-asymmetry sum rule can be further simplified into a more crude relation between the $CP$ asymmetries of $B^+\to K^+\pi^0$ and $B^0\to K^+\pi^-$ \cite{Gronau:1998ep}:
\begin{equation}\label{RoughRelation}
A_{CP}^{K^+\pi^0}-A_{CP}^{K^+\pi^-} \approx 0,
\end{equation}
which is clearly in contradiction to the latest world average of the $CP$-asymmetry difference between these two aforementioned processes \cite{Belle:2012dmz,BaBar:2012fgk,CDF:2014pzb,LHCb:2018pff,HFLAV:2019otj,LHCb:2020dpr},
\begin{equation}
\Delta A_{CP}(K\pi)\equiv A_{CP}^{K^+\pi^0}-A_{CP}^{K^+\pi^-}=0.115\pm0.014.
\end{equation}
This is basically a short version of the long-standing ``$K\pi$-puzzle''.

The $K\pi$-puzzle has gained constant theoretical attentions, especially those works on the exploration of the possibility of NP \cite{Buras:2003yc,Buras:2003dj,Buras:2004ub,Barger:2004hn,Baek:2004rp,Chang:2009wt,Fleischer:2017vrb,Beaudry:2017gtw,Bhattacharya:2021shk}.
Meanwhile, there are also works attempting to understand it within SM \cite{Bauer:2005kd,Cheng:2009cn,Cheng:2014rfa,Zhou:2016jkv,Kundu:2021emt}.
In this paper, we will try to examine the role of the isospin breaking effect --the $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effect-- in the $K\pi$-puzzle.

\section{the pollution of the $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effects to the $K\pi$-puzzle}

The basic idea is very simple.
Since there are $\pi^0$s in the final states of these $B\to K\pi$ decay processes, the isospin-breaking $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effect \cite{Gross:1979ur,Leutwyler:1996np,Feldmann:1998sh,Escribano:2020jdy,Kordov:2021eqx} takes place.
Although this effect seems to be negligible at first sight as it is small \cite{Gronau:2006eb} --perhaps that is why the $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effect has never been put on the table dealing with the $K\pi$-puzzle-- this kind of isospin-breaking effects could potentially affect $CP$ asymmetries more badly than expected \cite{Zhang:2020rol}.

Within the context of the $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effect, the $\pi^0$ meson can be expressed as an admixture of the isospin eigenstate $\pi_3$ and the mass eigenstates $\eta$ and $\eta^\prime$, $|\pi^0\rangle=|\pi_3\rangle+\epsilon|\eta\rangle+\epsilon^\prime|\eta^\prime\rangle$, where $\epsilon$ and $\epsilon^\prime$ are small parameters which account for the mixing between $\pi^0$ and $\eta$ and $\eta^\prime$, respectively \cite{Leutwyler:1996np,Feldmann:1998sh}.
The $CP$ asymmetry parameters of $B^{+}\to  K^{+}\pi^0$ and $B^{0}\to  K^{0}\pi^0$ acquire corrections because of the $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing.
For example, up to $\mathcal{O}({\epsilon^{(')}})$, the $CP$ asymmetry of $B^+\to K^+ \pi^0 $ is modified into
\footnote{An extra factor $\left(1-{A_{CP}^{K^+\pi_3 }}^2\right)$ in front of $\Delta_{\text{IB}}^{K^+\pi^0 }$ is omitted as it is numerically very close to one.}
\begin{equation}\label{ACPcorrection}
  A_{CP}^{ K^+ \pi^0}= A_{CP}^{ K^+\pi_3}+\Delta_{\text{IB}}^{K^+\pi^0 },
\end{equation}
where $A_{CP}^{K^+\pi_3 }$ is in fact the conventionally defined $CP$ asymmetry of $B^+\to K^+\pi^0$ without the $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effect, which takes the form $A_{CP}^{K^+\pi_3 }\equiv \frac{\left|\mathcal{A}_{B^-\to K^-\pi_3 }\right|^2-\left|\mathcal{A}_{B^+\to  K^+\pi_3}\right|^2}{\left|\mathcal{A}_{B^-\to  K^-\pi_3}\right|^2+\left|\mathcal{A}_{B^+\to  K^+\pi_3}\right|^2}$,
and $\Delta_{\text{IB}}^{ K^+\pi^0}$ represents the correction of the $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effect to the $CP$ asymmetry of $B^+\to K^+ \pi^0$, which, after some algebra, can be expressed as
\begin{equation}\label{DeltaIBKpm}
  \Delta_{\text{IB}}^{K^+\pi^0 }=\Re\left[\left(\frac{\mathcal{A}_{B^-\to K^-\eta}}{\mathcal{A}_{B^-\to K^-\pi^0 }}-\frac{\mathcal{A}_{B^+\to K^+\eta}}{\mathcal{A}_{B^+\to K^+\pi^0}}\right)e^{i\theta}\epsilon+\left(\frac{\mathcal{A}_{B^-\to K^-\eta^{\prime}}}{\mathcal{A}_{B^-\to K^-\pi^0}}-\frac{\mathcal{A}_{B^+\to K^+\eta^{\prime}}}{\mathcal{A}_{B^+\to K^+\pi^0}}\right)e^{i\theta'}\epsilon'\right].
\end{equation}
Note that two strong phases between the amplitudes of $B^\pm\to K^\pm\eta^{(\prime)}$ and $B^\pm\to K^\pm\pi^0$, $\theta$ and $\theta'$, are presented explicitly.
From Eq. (\ref{ACPcorrection}) and (\ref{DeltaIBKpm}) one can roughly see that the $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing correction term can not be simply neglected, provided that the $CP$ asymmetry parameter $A_{CP}^{ K^+ \pi^0}$ is about the same order with the mixing parameters $\epsilon$ and/or $\epsilon'$.
Similarly, there is also a correction term to the $CP$ asymmetry of $B^0\to K^0\pi^0$, which reads
\begin{equation}\label{DeltaIBK0}
  \Delta_{\text{IB}}^{ K^0\pi^0}=\Re\left[\left(\frac{\mathcal{A}_{\overline{B^0}\to \overline{K^0}\eta}}{\mathcal{A}_{\overline{B^0}\to \overline{K^0}\pi^0}}-\frac{\mathcal{A}_{B^0\to K^0\eta}}{\mathcal{A}_{B^0\to K^0\pi^0}}\right)e^{i\tilde{\theta}}\epsilon+\left(\frac{\mathcal{A}_{\overline{B^0}\to \overline{K^0}\eta^{\prime}}}{\mathcal{A}_{\overline{B^0}\to \overline{K^0}\pi^0}}-\frac{\mathcal{A}_{B^0\to K^0\eta^{\prime}}}{\mathcal{A}_{B^0\to K^0\pi^0}}\right)e^{i\tilde{\theta}'}\epsilon'\right].
\end{equation}

Since Eqs. (\ref{SumRule}) and (\ref{RoughRelation}) are obtained under the ignorance of the the $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effect, the $\pi^0$s in these two equations are in fact the isospin eigenstates $\pi_3$s.
Consequently, Eqs. (\ref{SumRule}) and (\ref{RoughRelation}) should be rewritten as
\begin{equation}\label{SumRuleNonPhysical}
A_{CP}^{K^+\pi^-}+\frac{\mathcal{B}(K^0\pi^+)}{\mathcal{B}(K^+\pi^-)}\frac{\tau_0}{\tau_+}A_{CP}^{K^0\pi^+} =\frac{2\mathcal{B}(K^+\pi_3)}{\mathcal{B}(K^+\pi^-)}\frac{\tau_0}{\tau_+}A_{CP}^{K^+\pi_3} +\frac{2\mathcal{B}(K^0\pi_3)}{\mathcal{B}(K^+\pi^-)}\frac{\tau_0}{\tau_+}A_{CP}^{K^0\pi_3},
\end{equation}
and
\begin{equation}\label{RoughRelationIB}
A_{CP}^{K^+\pi_3}-A_{CP}^{K^+\pi^-}\approx 0,
\end{equation}
respectively.
It should be pointed out that all the branching ratios and $CP$ asymmetries containing $\pi_3$ in the final state are strictly speaking not physical observables.
This however, can be fixed with the aid of Eq. (\ref{ACPcorrection}) by rewriting Eqs. (\ref{SumRuleNonPhysical})  and (\ref{RoughRelationIB}) in terms of the physical branching ratios and $CP$ asymmetries with $\pi^0$s containing in the final states.
The $CP$-asymmetry sum rule and the $CP$-asymmetry difference now read
\begin{equation}\label{SumRulePhysical}
A_{CP}^{K^+\pi^-}+\frac{\mathcal{B}(K^0\pi^+)}{\mathcal{B}(K^+\pi^-)}\frac{\tau_0}{\tau_+}A_{CP}^{K^0\pi^+} =\frac{2\mathcal{B}(K^+\pi^0)}{\mathcal{B}(K^+\pi^-)}\frac{\tau_0}{\tau_+}A_{CP}^{K^+\pi^0}+\frac{2\mathcal{B}(K^0\pi^0)}{\mathcal{B}(K^+\pi^-)}\frac{\tau_0}{\tau_+}A_{CP}^{K^0\pi^0}
-\Delta_{\text{IB}},
\end{equation}
and
\begin{equation}\label{RoughRelationIBPhysical}
A_{CP}^{K^+\pi^0}-A_{CP}^{K^+\pi^-}\approx\Delta_{\text{IB}},
\end{equation}
respectively, where $\Delta_{\text{IB}}$ accommodates the  $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing correction and takes the form
\footnote{Contributions from the branching ratios, which are proportional to $ A_{CP}^{K^{+,0}\pi^0} \epsilon^{(\prime)}$, have been neglected in $\Delta_{\text{IB}}$.}
\begin{eqnarray}
&&\Delta_{\text{IB}}=\frac{2\mathcal{B}(K^+\pi^0)}{\mathcal{B}(K^+\pi^-)}\frac{\tau_0}{\tau_+}\Delta_{\text{IB}}^{K^+\pi^0 } +\frac{2\mathcal{B}(K^0\pi^0)}{\mathcal{B}(K^+\pi^-)}\frac{\tau_0}{\tau_+}\Delta_{\text{IB}}^{ K^0\pi^0}.
\end{eqnarray}
One interesting behaviour of this modification is that although Eq. (\ref{RoughRelation}) relates only the $CP$ asymmetries of $B^0\to K^+\pi^-$ and $B^+\to K^+\pi^0$, the isospin-breaking correction term $\Delta_{\text{IB}}$ in Eq. (\ref{RoughRelationIBPhysical}), however, not only contains the contribution of the process $B^+\to K^+\pi^0$, but also contains that of the process $B^0\to K^0\pi^0$.
The latter turns out to be numerically even more important in $\Delta_{\text{IB}}$.

We are now in a position to estimate the impact of the the isospin-breaking correction term $\Delta_{\text{IB}}$ to the $K\pi$-puzzle.
Although the amplitudes in Eqs. (\ref{DeltaIBKpm}) and (\ref{DeltaIBK0}) can be calculated theoretically or extracted from data, the four relative strong phases between these amplitudes, $\theta$, $\theta'$, $\tilde{\theta}$, and $\tilde{\theta}'$, are non-perturbative, preventing us from an accurate prediction of $\Delta_{\text{IB}}$.
What we can do is to give an rough estimation of the possible range of $\Delta_{\text{IB}}$ by treating the four strong phases as free parameters varying from 0 to $2\pi$ independently.
Based on this strategy, with the amplitudes borrowed from Ref. \cite{Cheng:2014rfa}, $\Delta_{\text{IB}}$ is estimated to be
\begin{equation}\label{eq:rangeDeltaIB}
  \Delta_{\text{IB}}=(-0.37,+0.37)\times\epsilon+(-1.16,+1.16)\times\epsilon'.
\end{equation}

In TABLE \ref{tab:rangDeltaIB}, with different values of $\epsilon$ and $\epsilon'$ from different references, the corresponding allowed range of $\Delta_{\text{IB}}$ is calculated via Eq. (\ref{eq:rangeDeltaIB}).
Note that the mixing parameters $\epsilon$ and $\epsilon'$ take quite different values throughout the literature, resulting in quite different ranges of the $\Delta_{\text{IB}}$, as is presented in the last column of this table.
From TABLE \ref{tab:rangDeltaIB} one can see that the isospin-breaking-correction term $\Delta_{\text{IB}}$ can be as large as a few percent, which indicates that the influence of the $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effect to the $CP$-asymmetry sum rule of $B\to K\pi$ and the $CP$-asymmetry difference of the $B^+\to K^+\pi^0$ and $B^0\to K^+\pi^-$ can not be simply ignored.
It is even possible that the $K\pi$-puzzle can be explained by the $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effect {\it alone}.
Take the $CP$-asymmetry difference between $B^+\to K^+\pi^0$ and $B^0\to K^+\pi^-$ as example.
If we treat the $\Delta_{\text{IB}}$-term as an uncertainty, with the mixing parameters from Ref. \cite{Escribano:2020jdy}, the $CP$-asymmetry difference now should read as
\begin{equation}
A_{CP}^{K^+\pi_3}-A_{CP}^{K^+\pi^-}=0.115\pm0.014\pm0.041_{\pi^0\!\!-\!\eta\!-\!\eta'}=0.115\pm0.043,
\end{equation}
from which one can clearly see that the uncertainty caused by the $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effect can be even larger than the other experimental uncertainties combined.
Hence the significance of the nonzeroness of this $CP$ difference is considerably reduced to less that 3 standard deviations.

\begin{table}[h!]
  \begin{center}
    \caption{The $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing parameters quoted from different references and the corresponding range of $\Delta_{\text{IB}}$. Only the central values of $\epsilon$ and $\epsilon'$ are used when obtaining the range of $\Delta_{\text{IB}}$.}
    \label{tab:rangDeltaIB}
\begin{tabular}{c|c|c|c}
  \hline
  reference & $\epsilon (\%)$ & $\epsilon' (\%)$ & $\Delta_{\text{IB}} (\%$) \\
  \hline
  Kroll \cite{Kroll:2005sd} & $1.7\pm0.2$ & $0.4\pm0.1$ & $(-1.1,+1.1)$ \\
  Escribano \& Royo\cite{Escribano:2020jdy} & $0.1\pm0.9$ & $3.5\pm0.9$ & $(-4.1,+4.1)$ \\
  Benayoun et. al. \cite{Benayoun:2021ody} & $4.16\pm0.20$ & $1.05\pm0.05$ & $(-2.8,+2.8)$ \\
  \hline
\end{tabular}
  \end{center}
\end{table}

\section{Conclusion}
In conclusion, the contribution of the isospin-breaking $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effect to the $CP$ asymmetries of $B\to K\pi$ is investigated for the first time in this paper.
It is found that the $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effect pollutes the $K\pi$-puzzle more dramatically than the naive expectation.
This pollution is imbedded in the parameter $\Delta_{\text{IB}}$, which brings quite a large uncertainty to the $CP$-asymmetry sum rule of $B\to K\pi$ and the $CP$-asymmetry difference of the $B^+\to K^+\pi^0$ and $B^0\to K^+\pi^-$.
The analysis of this paper shows that it is even possible to explain the $K\pi$-puzzle by the $\pi^0\!\!-\!\eta\!-\!\eta'$ mixing effect {\it alone}.
Consequently, the implications of the $K\pi$-puzzle should be reconsidered.

\begin{acknowledgments}
This work was supported by National Natural Science Foundation of China under Contracts Nos. 11705081 and 12192261.
\end{acknowledgments}

\bibliography{zzhbib}
\end{document}